The pin - connected rigid rods AB and BC are inclined at theta = 30 when they are unloaded. When the force P is applied theta becomes 30.2. Determine the average normal strain developed in wire AC.

no title provided
Answer #1
Concepts and reason

Normal strain is a measure of longitudinal deformation that is elongation or contraction in the body. Normal strain is denoted by ε\varepsilon .

Normal stress can be calculated using known values of Young’s modulus and normal strain.

Normal stress is used in various real time applications.

Fundamentals

Force is a vector quantity. It is represented by a magnitude, direction, and point of application.

The direction of force can be either compressive or tensile.

The normal strain is calculated as follows:

Р
Р

From the figure, initial length is L1{L_1} and the final length is L2{L_2} .

The normal strain is,

ε=L2L1L1\varepsilon = \frac{{{L_2} - {L_1}}}{{{L_1}}}

Write the relation to calculate initial length of wire AC.

L1=2Lsinθ1{L_1} = 2L\sin {\theta _1}

Here, the length of the rod is L and the angle between the rods before load is applied is θ1{\theta _1} .

Substitute 0.6m0.6\,{\rm{m}} for L and 3030^\circ for θ1{\theta _1} .

L1=2×0.6×sin30=0.6m\begin{array}{c}\\{L_1} = 2 \times 0.6 \times \sin 30^\circ \\\\ = 0.6\,{\rm{m}}\\\end{array}

Write the relation to calculate final length of wire AC.

L2=2Lsinθ2{L_2} = 2L\sin {\theta _2}

Here, the angle between the rods after load is applied is θ2{\theta _2} .

Substitute 0.6m0.6\,{\rm{m}} for L and 30.230.2^\circ for θ1{\theta _1} .

L2=2×0.6×sin30.2=0.60362m\begin{array}{c}\\{L_2} = 2 \times 0.6 \times \sin 30.2^\circ \\\\ = 0.60362\,{\rm{m}}\\\end{array}

Write the relation to calculate the average normal strain developed in the wire AC.

ε=L2L1L1\varepsilon = \frac{{{L_2} - {L_1}}}{{{L_1}}}

Substitute 0.60362m0.60362\,{\rm{m}} for L2{L_2} and 0.6m0.6\,{\rm{m}} for L1{L_1} .

ε=0.603620.60.6=6.033×103\begin{array}{c}\\\varepsilon = \frac{{0.60362 - 0.6}}{{0.6}}\\\\ = 6.033 \times {10^{ - 3}}\,\\\end{array}

Ans:

The average normal strain developed in the wire AC is 6.033×1036.033 \times {10^{ - 3}}\, .